User avatar
sooted up Cyndi
4 Jan 2022 12:33 pm
User avatar
Water Cooler Poleece
Water Cooler Poleece
10,440 posts
Vegas » 04 Jan 2022, 1:22 pm » wrote: ↑ It looks like you are spot on, but there is a little more math that needs to be done for no other reason than to prove it mathematically.  I believe (though I may be wrong) that to find the real probabilities, there are 4 cases: 
  1. vaccine with adverse conditions without infection
  2. Vaccine with adverse conditions with infection
  3. vaccine without adverse conditions with infection
  4. Vaccine without adverse conditions without infection
Probabilities: 
  • Vaccine with adverse conditions = .2
  • Vaccine without adverse conditions = .8
  • Infected = 3/10000 = .03%
  • Not infected = 99.97%
Case 1) .2 x .9997 = 19.994%
Case 2) .2 x .0003 = .006%
Case 3) .8 x .03 = 2.4%
Case 4) .8 x .9997 = 79.98%

Case 4 has the highest probability followed by Case 1. However, these are only the probabilities with the vaccine. I would say that case 2 is the most applicable to your argument. Very small probability, so you are spot on.
Holy crap. :lol:  you a mathematician.. or do stats. :lol:  
Image
Updated 2 minutes ago
© 2012-2026 Liberal Forum

Search